# Canonical bases of tensor products of integrable highest weight modules arising from framed constructions

### v2

@@ 2.5.1. Canonical basis of $\mathbf{f}$
We refer [13, Chapter 14] and [11, Sections 10] for the definition of the canonical basis $\mathbf{B}$ of $\mathbf{f}$. We summarize some properties of $\mathbf{B}$ as follows.
@@
(almost orthogonal) for any $b,b^{\prime}\in\mathbf{B}$, we have $(b,b^{\prime})\in\delta_{b,b^{\prime}}+v^{-1}\mathbb{Z}[[v^{-1}]]$;
@@ 2.5.2. Canonical basis of $\Lambda_{\lambda}$
(almost orthogonal) for any $b,b^{\prime}\in\mathbf{B}(\Lambda_{\lambda})$, we have $(b,b^{\prime})_{\lambda}\in\delta_{b,b^{\prime}}+v^{-1}\mathbb{Z}[v^{-1}]$;
@@ 2.5.3. Canonical basis of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$
We refer [2] for the constructions of the canonical basis $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ of the tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ which is a generalization of [13, Chapter 24]. For convenience, we sketch the construction as follows.
For any $\nu\in\mathbb{N}[I]$, let $\{b^{*}\mid b\in\mathbf{B}_{\nu}\}$ be the basis of $\mathbf{f}_{\nu}$ dual to the canonical basis $\mathbf{B}_{\nu}$ under the inner product (,)(,)( , ) of $\mathbf{f}$. The $-\mathcal{R}$-matrix is the linear map $\Theta:\Lambda_{\xi}\otimes\Lambda_{\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ defined by
@@
Let $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})=\{b_{1}\otimes b_{2}\mid b_{1}\in\mathbf{B}(\Lambda_{\xi}),b_{2}\in\mathbf{B}(\Lambda_{\lambda})\}$. Since $\mathbf{B}(\Lambda_{\xi})$ and $\mathbf{B}(\Lambda_{\lambda})$ are almost orthogonal, by definitions, we have
$(b_{1}\otimes b_{2},b^{\prime}_{1}\otimes b^{\prime}_{2})_{\xi,\lambda}=(b_{1},b^{\prime}_{1})_{\xi}(b_{2},b^{\prime}_{2})_{\lambda}\in\delta_{b_{1}\otimes b_{2},b^{\prime}_{1}\otimes b^{\prime}_{2}}+v^{-1}\mathbb{Z}[v^{-1}]$
for any $b_{1}\otimes b_{2},b^{\prime}_{1}\otimes b^{\prime}_{2}\in\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$, and so $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$ is an almost orthogonal basis of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$. Let $\mathcal{L}_{\xi,\lambda}$ be the $\mathbb{Z}[v^{-1}]$-submodule of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ generated by $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$.
@@ 3. Framed construction of the tensor product
We fix two dominant weights $\xi,\lambda\in X$ with respect to the root datum $(Y,X,\langle,\rangle,...)$.
@@ 3.1. Framed Cartan datum and root datum
For the Cartan datum $(I,\cdot)$, we define its framed Cartan datum to be $(\tilde{I},\cdot)$, where $\tilde{I}=I\sqcup I^{\prime}=I\sqcup\{i^{\prime}\mid i\in I\}$, and the symmetric bilinear form $\mathbb{Z}[\tilde{I}]\times\mathbb{Z}[\tilde{I}]\rightarrow\mathbb{Z}$ is extended from the bilinear form in $(I,\cdot)$ such that
@@
Then $(\tilde{I},\cdot)$ is a symmetric Cartan datum. For the root datum $(Y,X,\langle,\rangle,...)$, suppose that $Y=Y^{0}\oplus\mathbb{Z}[I]$ and $X\cong\mathrm{Hom}_{\mathbb{Z}}(Y^{0},\mathbb{Z})\oplus\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}[I],\mathbb{Z})$, we define its framed root datum to be $(\tilde{Y},\tilde{X},\langle,\rangle,...)$, where $\tilde{Y}=Y\oplus\mathbb{Z}[I^{\prime}],\ \tilde{X}=X\oplus\mathbb{Z}[\{i^{\prime*}\mid i\in I\}]$ with the perfect bilinear pairing $\tilde{Y}\times\tilde{X}\rightarrow\mathbb{Z}$ extended from the pairing in $(\tilde{Y},\tilde{X},\langle,\rangle,...)$, and the embeddings $\tilde{I}\hookrightarrow\tilde{Y},\tilde{I}\hookrightarrow\tilde{X}$ are extended from the embeddings in $(\tilde{Y},\tilde{X},\langle,\rangle,...)$ such that
@@
Then $(\tilde{Y},\tilde{X},\langle,\rangle,...)$ is a $\tilde{Y}$-regular root datum of type $(\tilde{I},\cdot)$. The embeddings $\tilde{I}\hookrightarrow\tilde{Y}$ and $\tilde{I}\hookrightarrow\tilde{X}$ induce homomorphisms $\mathbb{Z}[\tilde{I}]\rightarrow\tilde{Y}$ and $\mathbb{Z}[\tilde{I}]\rightarrow\tilde{X}$, we shall often denote the images of $\nu\in\mathbb{Z}[\tilde{I}]$ under either of these homomorphisms by the same notation $\nu$.
For the framed Cartan datum $(\tilde{I},\cdot)$, we denote by $\tilde{\mathbf{f}}$ the corresponding algebra defined in 2.1. For the framed root datum $(\tilde{Y},\tilde{X},\langle,\rangle,...)$, we denote by $\tilde{\mathbf{U}}$ the corresponding algebra defined in 2.2. It is clear that the natural embedding $I\hookrightarrow\tilde{I}$ induces subalgebra embeddings $\mathbf{f}\hookrightarrow\tilde{\mathbf{f}}$ and $\mathbf{U}\hookrightarrow\tilde{\mathbf{U}}$.
@@ 3.2. Framed construction of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$
Note that it is well-defined, since $i^{\prime}\cdot j^{\prime}=0$ and $\theta_{i^{\prime}}\theta_{j^{\prime}}=\theta_{j^{\prime}}\theta_{i^{\prime}}$ for any $i\not=j$ in $I$ by the quantum Serre relations. We can identify $\lambda$ with the sequence $(\langle i,\lambda\rangle i^{\prime})_{i\in I}\in\mathcal{V}_{|\theta_{\lambda}|}$ in arbitrary order such that $\theta_{\lambda}$ is accordance with the notation $\theta_{(\langle i,\lambda\rangle i^{\prime})_{i\in I}}$.
@@
In particular, $\xi\odot\lambda\in\tilde{X}$ is dominant with respect to the framed root datum $(\tilde{Y},\tilde{X},\langle,\rangle,...)$. We denote by $\tilde{\Lambda}_{\xi\odot\lambda}$ the simple integrable highest weight module of $\tilde{\mathbf{U}}$ defined in 2.3. Then $\tilde{M}_{\xi\odot\lambda}$ and $\tilde{\Lambda}_{\xi\odot\lambda}$ are $\mathbf{U}$-modules via the natural embedding $\mathbf{U}\hookrightarrow\tilde{\mathbf{U}}$.
@@
It is well-known that any submodule of a weight module is also a weight module, and so we only need to prove that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is a $\mathbf{U}$-submodule of $\tilde{M}_{\xi\odot\lambda}$. It is clear that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is stable under the actions of $F_{i}$ and $K_{\mu}$ for any $i\in I$ and $\mu\in Y$. We need to prove that it is stable under the actions of $E_{i}$ for any $i\in I$. For any homogeneous $x,y\in\mathbf{f}$, we make an induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that $E_{i}x\theta_{\lambda}y\in\mathbf{f}\theta_{\lambda}\mathbf{f}$.
@@
$E_{i}\theta_{\lambda}y=E_{i}\prod_{i\in I}F_{i^{\prime}}^{(\langle i,\lambda\rangle)}y=\prod_{i\in I}F_{i^{\prime}}^{(\langle i,\lambda\rangle)}E_{i}y=\theta_{\lambda}E_{i}y.$
We make an induction on $\mathrm{tr}\,|y|\in\mathbb{N}$ to prove that $E_{i}y\in\mathbf{f}$ which implies that $E_{i}\theta_{\lambda}y\in\mathbf{f}\theta_{\lambda}\mathbf{f}$.
[…]
$E_{i}y=E_{i}\theta_{j}y^{\prime}=E_{i}F_{j}y^{\prime}=F_{j}E_{i}y^{\prime}+\delta_{i,j}[\langle i,\xi\odot\lambda-|y^{\prime}|\rangle]y^{\prime}\in F_{j}\mathbf{f}+\mathbf{f}\subset\mathbf{f},$
as desired.
@@
For any $i\in I$, we make an induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that $(E_{i}x)\theta_{\lambda}=E_{i}(x\theta_{\lambda})$.
@@
and so $\cdot\theta_{\lambda}$ induces a well-defined $\mathbf{U}$-module homomorphism $\psi:\Lambda_{\lambda}\rightarrow\Lambda_{0,\lambda}$. It is clear that $\psi$ is surjective, and $\Lambda_{0,\lambda}\not=0$. Since $\Lambda_{\lambda}$ is a simple integrable highest weight module, $\psi$ is an isomorphism. ∎
@@
[…]
@@
For any homogeneous $x,y\in\mathbf{f}$, we calculate the image $\varphi^{\prime\prime}(x\theta_{\lambda}y)$ as follows. Suppose that $r(x)=\sum x_{1}\otimes x_{2},r(y)=\sum y_{1}\otimes y_{2}$, where $x_{1},x_{2},y_{1},y_{2}\in\mathbf{f}$ are homogeneous, by definitions, we have
@@
Any surjective linear maps from $\Lambda_{\xi,\lambda}$ to $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ preserving weight subspaces are linear isomorphism.
@@
For any $i\in I$, we make an induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that $\varphi^{\prime}(E_{i}x\theta_{\lambda}y)=E_{i}\varphi^{\prime}(x\theta_{\lambda}y)$.
$\bullet$ If $\mathrm{tr}\,|x|=0$, recall that we have proved that $E_{i}y\in\mathbf{f}$ in the proof of Lemma 3.1, and so we can suppose that $E_{i}y=\sum c_{y^{\prime}}y^{\prime}$, where $y^{\prime}\in\mathbf{f}$ are homogeneous and $c_{y^{\prime}}\in\mathbb{Q}(v)$. By Corollary 3.6, we have
$\varphi^{\prime}(E_{i}\theta_{\lambda}y)=$ $\varphi^{\prime}(E_{i}\prod_{i\in I}F_{i^{\prime}}^{(\langle i,\lambda\rangle)}y)=\varphi^{\prime}(\prod_{i\in I}F_{i^{\prime}}^{(\langle i,\lambda\rangle)}E_{i}y)=\varphi^{\prime}(\theta_{\lambda}\sum c_{y^{\prime}}y^{\prime})$
$=$ $\sum c_{y^{\prime}}{y^{\prime}}^{-}\eta_{\xi}\otimes\eta_{\lambda}=E_{i}y^{-}\eta_{\xi}\otimes\eta_{\lambda}+K_{i}y^{-}\eta_{\xi}\otimes E_{i}\eta_{\lambda}=E_{i}\varphi^{\prime}(\theta_{\lambda}y).$
@@
as desired. Hence the map $\varphi^{\prime}$ is a $\mathbf{U}$-module homomorphism. Secondly, we prove that $\varphi^{\prime}(\mathbf{f}\theta_{\lambda}\mathbf{f}\cap\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1})=0$. It suffices to prove that
@@
$\tau(p\otimes\mathrm{Id})r(\sum_{i\in I}x_{i}\theta^{(\langle i,\xi\rangle+1)})$
@@
$\varphi^{\prime\prime}(\sum_{i\in I}x_{i}\theta_{i}^{\langle i,\xi\rangle+1})=\sum_{i\in I}\sum x_{i2}\theta_{i}^{(\langle i,\xi\rangle+1)}\otimes\pi_{0\odot\lambda}(x_{i1})\in\sum_{i\in I}\mathbf{f}\theta_{i}^{\langle i,\xi\rangle+1}\otimes\Lambda_{0,\lambda}.$
Hence $\varphi^{\prime}$ induces a well-defined $\mathbf{U}$-module homomorphism $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$. Thirdly, we prove that $\varphi$ is an isomorphism. Notice that the vector space $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ is spanned by $\{y^{-}\eta_{\xi}\otimes x^{-}\eta_{\lambda}\mid x,y\in\mathbf{f}\ \textrm{arehomogeneous}\}$. We make an induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that $y^{-}\eta_{\xi}\otimes x^{-}\eta_{\lambda}\in\mathrm{Im}\,\varphi$.
@@
It is clear that the bar-involution of $\tilde{\mathbf{f}}$ can be restricted to be a bar-involution of the subspace $\mathbf{f}\theta_{\lambda}\mathbf{f}$, and the bar-involution $\tilde{\Lambda}_{\xi\odot\lambda}$ can be restricted to be a bar-involution of the submodule $\Lambda_{\xi,\lambda}$ such that $\pi_{\xi\odot\lambda}(\overline{x})=\overline{\pi_{\xi\odot\lambda}(x)}$ for any $x\in\mathbf{f}\theta_{\lambda}\mathbf{f}$.
@@
It suffices to prove that $\varphi^{\prime}(\overline{x})=\Psi\varphi^{\prime}(x)$ for any $x\in\mathbf{f}\theta_{\lambda}\mathbf{f}$. Notice that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is spanned by $\{\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}\mid\boldsymbol{\nu}_{1}\in\mathcal{V}_{\nu_{1}},\boldsymbol{\nu}_{2}\in\mathcal{V}_{\nu_{2}},\nu_{1},\nu_{2}\in\mathbb{N}[I]\}$. We make an induction on $\mathrm{tr}\,\nu_{1}\in\mathbb{N}$ to prove that $\varphi^{\prime}(\overline{\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}})=\Psi\varphi^{\prime}(\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}})$.
@@
By Proposition 2.3, there is a unique admissible (with respect to the $\tilde{\mathbf{U}}$-action) symmetric bilinear inner product $(,)_{\xi\odot\lambda}:\tilde{\Lambda}_{\xi\odot\lambda}\times\tilde{\Lambda}_{\xi\odot\lambda}\rightarrow\mathbb{Q}(v)$ such that $(\eta_{\xi\odot\lambda},\eta_{\xi\odot\lambda})_{\xi\odot\lambda}=1$.

### v3

@@ 2.5.1. Canonical basis of $\mathbf{f}$
We refer to [13, Chapter 14] and [11, Sections 10] for the definition of the canonical basis $\mathbf{B}$ of $\mathbf{f}$. We summarize some properties of $\mathbf{B}$ as follows.
@@
(almost orthonormal) for any $b,b^{\prime}\in\mathbf{B}$, we have $(b,b^{\prime})\in\delta_{b,b^{\prime}}+v^{-1}\mathbb{Z}[[v^{-1}]]$;
@@ 2.5.2. Canonical basis of $\Lambda_{\lambda}$
(almost orthonormal) for any $b,b^{\prime}\in\mathbf{B}(\Lambda_{\lambda})$, we have $(b,b^{\prime})_{\lambda}\in\delta_{b,b^{\prime}}+v^{-1}\mathbb{Z}[v^{-1}]$;
@@ 2.5.3. Canonical basis of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$
We refer to [2] for the constructions of the canonical basis $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ of the tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ which is a generalization of [13, Chapter 24]. For convenience, we sketch the construction as follows.
For any $\nu\in\mathbb{N}[I]$, let $\{b^{*}\mid b\in\mathbf{B}_{\nu}\}$ be the basis of $\mathbf{f}_{\nu}$ dual to the canonical basis $\mathbf{B}_{\nu}$ under the bilinear form $(,)$ of $\mathbf{f}$. The quasi-$\mathcal{R}$-matrix is the linear map $\Theta:\Lambda_{\xi}\otimes\Lambda_{\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ defined by
@@
Let $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})=\{b_{1}\otimes b_{2}\mid b_{1}\in\mathbf{B}(\Lambda_{\xi}),b_{2}\in\mathbf{B}(\Lambda_{\lambda})\}$. Since $\mathbf{B}(\Lambda_{\xi})$ and $\mathbf{B}(\Lambda_{\lambda})$ are almost orthonormal, by definitions, we have
$(b_{1}\otimes b_{2},b^{\prime}_{1}\otimes b^{\prime}_{2})_{\xi,\lambda}=(b_{1},b^{\prime}_{1})_{\xi}(b_{2},b^{\prime}_{2})_{\lambda}\in\delta_{b_{1}\otimes b_{2},b^{\prime}_{1}\otimes b^{\prime}_{2}}+v^{-1}\mathbb{Z}[v^{-1}]\subset\delta_{b_{1}\otimes b_{2},b^{\prime}_{1}\otimes b^{\prime}_{2}}+v^{-1}\mathbb{Z}[[v^{-1}]]$
for any $b_{1}\otimes b_{2},b^{\prime}_{1}\otimes b^{\prime}_{2}\in\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$, and so $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$ is an almost orthonormal basis of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$. Let $\mathcal{L}_{\xi,\lambda}$ be the $\mathbb{Z}[v^{-1}]$-submodule of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ generated by $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$.
@@ 3. Framed construction of the tensor product
[…]
@@ 3.1. Framed Cartan datum and root datum
[…]
@@
[…]
@@
[…]
@@ 3.2. Framed construction of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$
Note that it is well-defined, since $i^{\prime}\cdot j^{\prime}=0$ and $\theta_{i^{\prime}}\theta_{j^{\prime}}=\theta_{j^{\prime}}\theta_{i^{\prime}}$ for any $i\not=j$ in $I$ by the quantum Serre relations. Consider the sequence $(\langle i,\lambda\rangle i^{\prime})_{i\in I}\in\mathcal{V}_{|\theta_{\lambda}|}$ in arbitrary order, then $\theta_{\lambda}$ is accordance with the notation $\theta_{(\langle i,\lambda\rangle i^{\prime})_{i\in I}}$ defined in §2.1.
@@
[…]
@@
It is well-known that any submodule of a weight module is also a weight module. It is clear that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is stable under the actions of $F_{i}$ and $K_{\mu}$ for any $i\in I$ and $\mu\in Y$. We need to prove that it is stable under the actions of $E_{i}$ for any $i\in I$. For any homogeneous $x,y\in\mathbf{f}$, we use induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that $E_{i}x\theta_{\lambda}y\in\mathbf{f}\theta_{\lambda}\mathbf{f}$.
@@
$E_{i}\theta_{\lambda}y=E_{i}\prod_{j\in I}F_{j^{\prime}}^{(\langle j,\lambda\rangle)}y=\prod_{j\in I}F_{j^{\prime}}^{(\langle j,\lambda\rangle)}E_{i}y=\theta_{\lambda}E_{i}y.$
Since $\mathbf{f}$ has a $\mathbf{U}$-module structure, it is claear that $E_{i}y\in\mathbf{f}$, and so $E_{i}\theta_{\lambda}y\in\mathbf{f}\theta_{\lambda}\mathbf{f}$.
@@
For any $i\in I$, we use induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that $(E_{i}x)\theta_{\lambda}=E_{i}(x\theta_{\lambda})$.
@@
and so $\cdot\theta_{\lambda}$ induces a well-defined $\mathbf{U}$-module homomorphism $\psi:\Lambda_{\lambda}\rightarrow\Lambda_{0,\lambda}$. It is clear that $\psi$ is surjective, and $\Lambda_{0,\lambda}\not=0$. Since $\Lambda_{\lambda}$ is an integrable highest weight module, $\psi$ is an isomorphism. ∎
@@
$\varphi^{\prime\prime}:$ $\mathbf{f}\theta_{\lambda}\mathbf{f}\hookrightarrow\tilde{\mathbf{f}}\xrightarrow{r}\tilde{\mathbf{f}}\otimes\tilde{\mathbf{f}}\xrightarrow{p\otimes\mathrm{Id}}\bigoplus_{\nu\in\mathbb{N}[I]}\tilde{\mathbf{f}}_{\nu+|\theta_{\lambda}|}\otimes\tilde{\mathbf{f}}\xrightarrow{\tau}\tilde{\mathbf{f}}\otimes\bigoplus_{\nu\in\mathbb{N}[I]}\tilde{\mathbf{f}}_{\nu+|\theta_{\lambda}|}$
@@
For any homogeneous $x,y\in\mathbf{f}$, we calculate the image $\varphi^{\prime\prime}(x\theta_{\lambda}y)$ as follows. Suppose that $r(x)=\sum x_{1}\otimes x_{2},r(y)=\sum y_{1}\otimes y_{2}$, where $x_{1},x_{2},y_{1},y_{2}\in\mathbf{f}$ are homogeneous. By definitions, we have
@@
Let $\mathbf{U}^{0}$ be the subalgebra $\mathbf{U}$ generated by $K_{\mu}$ for any $\mu\in Y$. Then any surjective $\mathbf{U}^{0}$-module homomorphisms from $\Lambda_{\xi,\lambda}$ to $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ are $\mathbf{U}^{0}$-module isomorphisms.
@@
For any $i\in I$, we use induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that $\varphi^{\prime}(E_{i}x\theta_{\lambda}y)=E_{i}\varphi^{\prime}(x\theta_{\lambda}y)$.
$\bullet$ If $\mathrm{tr}\,|x|=0$, recall that we have proved that $E_{i}y\in\mathbf{f}$ in the proof of Lemma 3.1. By Corollary 3.6, we have
$\varphi^{\prime}(E_{i}\theta_{\lambda}y)=$ $\varphi^{\prime}(E_{i}\prod_{i\in I}F_{i^{\prime}}^{(\langle i,\lambda\rangle)}y)=\varphi^{\prime}(\prod_{i\in I}F_{i^{\prime}}^{(\langle i,\lambda\rangle)}E_{i}y)=\varphi^{\prime}(\theta_{\lambda}(E_{i}y))=(E_{i}y)^{-}\eta_{\xi}\otimes\eta_{\lambda}.$
Since $E_{i}1=0$ in $\tilde{M}_{\xi\odot\lambda}$ and $E_{i}\eta_{\xi}=0$ in $\Lambda_{\xi}$, by [13, Proposition 3.1.6], we have
$E_{i}y=E_{i}y^{-}1=\frac{K_{i}{(_{i}r(y))}^{-}1-(r_{i}(y))^{-}K_{-i}1}{v-v^{-1}}=\frac{v^{\langle i,\xi\odot\lambda-|y|\rangle+2}{{}_{i}r(y)}-v^{\langle-i,\xi\odot\lambda\rangle}r_{i}(y)}{v-v^{-1}},$
$E_{i}y^{-}\eta_{\xi}=\frac{K_{i}{(_{i}r(y))}^{-}-(r_{i}(y))^{-}K_{-i}}{v-v^{-1}}\eta_{\xi}=\frac{v^{\langle i,\xi-|y|\rangle+2}({{}_{i}r(y)})^{-}-v^{\langle-i,\xi\rangle}(r_{i}(y))^{-}}{v-v^{-1}}\eta_{\xi}.$
Since $\langle i,\xi\odot\lambda\rangle=\langle i,\xi\rangle$, we have $(E_{i}y)^{-}\eta_{\xi}=E_{i}y^{-}\eta_{\xi}$, and so
$\varphi^{\prime}(E_{i}\theta_{\lambda}y)=(E_{i}y)^{-}\eta_{\xi}\otimes\eta_{\lambda}=E_{i}y^{-}\eta_{\xi}\otimes\eta_{\lambda}=E_{i}(y^{-}\eta_{\xi}\otimes\eta_{\lambda})=E_{i}\varphi^{\prime}(\theta_{\lambda}y).$
@@
as desired. Hence the map $\varphi^{\prime}$ is a $\mathbf{U}$-module homomorphism.
Secondly, we prove that $\varphi^{\prime}(\mathbf{f}\theta_{\lambda}\mathbf{f}\cap\sum_{i\in I}\tilde{\mathbf{f}}\theta_{i}^{\langle i,\xi\rangle+1})=0$. It suffices to prove that
@@
$\tau(p\otimes\mathrm{Id})r(\sum_{i\in I}x_{i}\theta_{i}^{(\langle i,\xi\rangle+1)})$
@@
$\varphi^{\prime\prime}(\sum_{i\in I}x_{i}\theta_{i}^{(\langle i,\xi\rangle+1)})=\sum_{i\in I}\sum x_{i2}\theta_{i}^{(\langle i,\xi\rangle+1)}\otimes\pi_{0\odot\lambda}(x_{i1})\in\sum_{i\in I}\mathbf{f}\theta_{i}^{\langle i,\xi\rangle+1}\otimes\Lambda_{0,\lambda}.$
Hence $\varphi^{\prime}$ induces a well-defined $\mathbf{U}$-module homomorphism $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$.
Thirdly, we prove that $\varphi$ is an isomorphism. Notice that the vector space $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ is spanned by $\{y^{-}\eta_{\xi}\otimes x^{-}\eta_{\lambda}\mid x,y\in\mathbf{f}\ \textrm{are homogeneous}\}$. We use induction on $\mathrm{tr}\,|x|\in\mathbb{N}$ to prove that $y^{-}\eta_{\xi}\otimes x^{-}\eta_{\lambda}\in\mathrm{Im}\,\varphi$.
@@
We remark that Theorem 3.8 still hold for any symmetrizable Cartan datum.
It is clear that the bar-involution of $\tilde{\mathbf{f}}$ can be restricted to be a bar-involution of the subspace $\mathbf{f}\theta_{\lambda}\mathbf{f}$, and the bar-involution of $\tilde{\Lambda}_{\xi\odot\lambda}$ can be restricted to be a bar-involution of the submodule $\Lambda_{\xi,\lambda}$ such that $\pi_{\xi\odot\lambda}(\overline{x})=\overline{\pi_{\xi\odot\lambda}(x)}$ for any $x\in\mathbf{f}\theta_{\lambda}\mathbf{f}$.
@@
It suffices to prove that $\varphi^{\prime}(\overline{x})=\Psi\varphi^{\prime}(x)$ for any $x\in\mathbf{f}\theta_{\lambda}\mathbf{f}$. Notice that $\mathbf{f}\theta_{\lambda}\mathbf{f}$ is spanned by $\{\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}\mid\boldsymbol{\nu}_{1}\in\mathcal{V}_{\nu_{1}},\boldsymbol{\nu}_{2}\in\mathcal{V}_{\nu_{2}},\nu_{1},\nu_{2}\in\mathbb{N}[I]\}$. We use induction on $\mathrm{tr}\,\nu_{1}\in\mathbb{N}$ to prove that $\varphi^{\prime}(\overline{\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}}})=\Psi\varphi^{\prime}(\theta_{\boldsymbol{\nu}_{1}}\theta_{\lambda}\theta_{\boldsymbol{\nu}_{2}})$.
@@
By Proposition 2.3, there is a unique admissible (with respect to the $\tilde{\mathbf{U}}-$action) symmetric bilinear form $(,)_{\xi\odot\lambda}:\tilde{\Lambda}_{\xi\odot\lambda}\times\tilde{\Lambda}_{\xi\odot\lambda}\rightarrow\mathbb{Q}(v)$ such that $(\eta_{\xi\odot\lambda},\eta_{\xi\odot\lambda})_{\xi\odot\lambda}=1$.